# SAMM: Sharded Automated Market Makers

### v1

@@ 8.5 Simulation of Trading Data
To evaluate SAMM in a real trading environment, we simulate its behavior using a performance profile based on our performance evaluation and workload from publicly available trading data of the Ethereum Uniswap v2 USDC-ETH pool, from Ethereum block $12,000,000$ to $19,500,000$ (from 2021-03-08 to 2024-03-23, about 3 years). We simulate Uniswap v2 and SAMM using $1$ to $32$ shards. We test three different values of $c$, namely $0.003,0.005$, and $0.01$, and choose other parameters through an optimization problem (Appendix I). In particular, there are no arbitrageurs, unlike the theoretical model.
We run each simulation instance as follows. We randomly select a point in Uniswap v2’s history as the starting point. For SAMM with $n$ shards, we evenly distribute the liquidity of the Uniswap v2 contract at that point among the $n$ shards to establish the initial state of SAMM. We then simulate the real trades from that time point onward. Each trade in the real data is simulated as a trading demand with the required amount of tokens, matching the output amount and tokens of the actual trade. To minimize costs, the trader selects the shard offering the lowest price and may split the trade into several smaller trades if this reduces her costs, disregarding gas fees. We assume both the Uniswap and SAMM pools operate at maximal throughput, consistent with the throughput in our performance evaluation. We count trade splits, the number of transactions in different SAMM shards, liquidity provider revenue, and trader costs over 1 second, following a 1-second warm-up period. We repeat each simulation 100 times and calculate averages. Note that $1$ second in our simulation corresponds to several hours in the real world. We get historical prices of ETH to USDC from the Yahoo Finance webpage [^2] https://finance.yahoo.com/quote/ETH-USD/history/ and use it as the external market price. We calculate the revenue for liquidity providers and costs for traders in USDC, converting ETH amounts at the real-time price.
We first analyze the incentives of liquidity providers and traders in SAMM. Figure 4(a) shows the average ratio of traders’ costs in SAMM compared to the costs in the external market, considering different numbers of shards and varying values of $c$. We observe that with larger $c$, the cost for traders increases, as expected. For a fixed $c$, the cost for traders decreases as the number of shards increases, aligning with SAMM’s $c$-smaller-better property, where more shards result in less liquidity in each shard. Additionally, we compare the cost for traders in Uniswap v2, depicted by the black line. In all cases, the cost of SAMM is either smaller ($c=0.003$ with no less than 2 shards) or slightly larger than that of Uniswap v2, differing by less than $1\%$ with $c=0.01$ and $0.3\%$ with $c=0.005$, when there are at least 2 shards. We aggregate trading fees across all shards to evaluate the revenue of liquidity providers. Figure 4(b) shows SAMM consistently generates higher revenue than Uniswap v2, with $c=0.01$ and 7 shards yielding over 15 times the revenue. Initially, revenue in SAMM rises with more shards due to increased throughput but with even more shards it declines as trading fees per trade decrease in smaller shards, as explained in Lemma 7.2.1.
@@
We monitor the distribution of trades across each shard in the SAMM system. Figure 5 illustrates the number of trades executed in each shard compared to the average. Error bars show the range of relative trade numbers in each shard compared to the average. We see where the difference from average is always under $5\%$. Additionally, we analyze the proportion of trades that are split into multiple transactions within SAMM. Figure 6 shows that with $c=0.01$ trade splits are infrequent, occurring in less than $8\%$ of trades, even with 32 shards. Furthermore, the total number of trade splits results in less than a $45\%$ increase in overall trade volume, a relatively minor increment given the five-fold throughput improvement. When the number of shards is larger or $c$ is smaller, the proportion of trade splits is higher.
@@
In summary, our simulation demonstrates that SAMM significantly outperforms Uniswap v2 in terms of liquidity provider revenue while maintaining comparable costs for traders. This increase without major cost hikes is due to SAMM’s enhanced throughput, allowing for more trades and higher total trading fees. The simulation also confirms that our theoretical predictions align with the actual trading behavior. Additionally, the results suggest that SAMM requires a larger $c$ for more shards to prevent trade splits. Although larger $c$ leads to larger trader costs, the cost remains acceptable due to the increased number of shards.
9 Conclusion
@@
We present SAMM, a scalable AMM. The key enabler of SAMM is the design of a trading fee mechanism that incentivizes parallel operations. We analyze trader and liquidity provider behaviors as a game, showing that parallel operations are the best response, and validate by simulation with real trade traces. We implement and deploy SAMM in a local Sui testnet, demonstrating more than 5x throughput improvement, up to the underlying system’s limits. Our results indicate that reducing serial bottlenecks of independent contracts should be a focus of smart-contract platforms to allow for AMM scaling (See Appendix H). Meanwhile, SAMM can be directly deployed to scale AMMs on existing platforms, for direct use and as part of the DeFi eco-system.
@@ Appendix A Growing Demand for AMM
We analyze the demand for AMMs by calculating the number of trades per second on the prominent Uniswap versions 1-3, from its deployment in November 2018 to March 2024. We use data from Dune [^3] https://dune.com/, a comprehensive database for blockchain data. Figure 7 illustrates the exponentially increasing demand, from $0.78$ average trades per second in 2020 to $9.54$ in 2024. On the Ethereum blockchain, most Uniswap transactions are executed through versions 2 and 3, incurring gas costs of $152,809$ and $184,523$ respectively. Given that the average total gas per block is $15,000,000$ and the block interval is $12$ seconds, Ethereum can facilitate up to $8.18$ trades per second for v2 and $6.77$ trades per second for v3. However, in March 2024, the average monthly trades on Uniswap reached $13.9$ per second, nearly doubling Ethereum’s processing capacity. Therefore, most demand of Uniswap is processed through off-chain solutions [2].
@@ Appendix C CPMM Does Not Satisfy Either property
{restatable}
thmthmcpmmlimitation For any value of $10<c<$ 1, the CPMM cost function $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$ does not satisfy either the $c$-Non-Splitting or the $c$-smaller-better properties.
Proof.
@@
Appendix D Proof of Proposition 5.3
  • Proposition 5.3 (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
The $c$-smaller-better property considers two pools $i$ and $j$, assuming their reported prices are identical. Denote the inverse of this price by $c^{\textit{AB}}$, so $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}=\frac{1}{c^{\textit{AB}}}>0$. The property requirement (Equation 5.1.2) becomes
@@
If the polynomial value value is greater than the bound, $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}>r_{\max}$, then from Equations 49, and 10, the gross amount becomes
@@
Combining the constraints we just found (Equations 32 and 33) into the gross amount expression (Equations 9 and 10) we have
@@
From Equations 28 and 33, we have
@@
Appendix E Proof of Theorem 5.3
  • Theorem 5.3 (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
Then we can expand the trading fee function and the gross amount function as:
@@
$\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{B}}{(1-c)^{3}\left(R^{A}\right)^{2}}\leq 0\,\,.$ (51)
The above inequation is equal to
@@
The above condition is sufficient for $c$-Non-Splitting property.
@@
Therefore, Item 1 is sufficient for Equation 52, which indicates the $c$-Non-Splitting property holds under Item 1.
@@
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}=-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}\,\,.$
@@
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}>-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-cR^{A})^{2}}\,\,.$ (55)
@@
The above equation is equal to
@@ Appendix F CPMM Equilibrium
In CPMM, the gross amount is linear to the net amount. Since traders can suffer less from slippage by splitting a trade, the gross amount of splitting a trade is less than the gross amount of trading the same amount at one time. {restatable}thmtheoremuniswapsplitting In $\Gamma_{n}(\textit{tf}_{\textit{CPMM}})$, the following strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}})$ is the only dominant strategy for the BA trader, where
@@
Proof.
We start by calculating the best response for the BA trader. Since the utility $U^{BA}(\textbf{R},b^{\textit{BA}},\pi^{BA})$ is a linear combination of the revenue over actions $U^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ (Equation 15), we can first calculate the optimal action for the BA trader. Combining the CPMM gross amount (Equation 4.3) and the revenue function (Equation 6.1.4), we obtain
@@
We first consider the case of two pools and then extend it to the general case. We define the function $f(\cdot,\cdot)$ which is proportional to the utility of trader in $\textit{pool}_{i}$ and $\textit{pool}_{j}$:
$f(b^{\textit{BA}}_{i},b^{\textit{BA}}_{j})\mathrel{\mathop{:}}=-\left(\frac{R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}+\frac{R^{A}_{j}b^{\textit{BA}}_{j}}{R^{A}_{j}-b^{\textit{BA}}_{j}}\right)\,\,,$
. Denote by $z\mathrel{\mathop{:}}=b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j}\leq b^{\textit{BA}}$. Then we have
@@
$\tau^{BA}(\textbf{R},b^{\textit{BA}})=\begin{cases}1,&\text{if }a^{\textit{BA}}=\left(\frac{R^{A}_{1}}{\sum R^{A}_{i}},\cdots,\frac{R^{A}_{n}}{\sum R^{A}_{i}}\right)\\ 0,&\text{Otherwise.}\end{cases}\qed$ (62)
Appendix G Equilibrium Proof Details
We provide the details for the proofs outlined in Section 7.
G.1 Proof of Lemma 7.1.1
  • Lemma 7.1.1 (restated). In $\Gamma_{n}(tf_{\textit{SAMM}})$, a trader wants to get $b^{\textit{BA}}$ token A when the system state is R. Then for the action which obtains all $b^{\textit{BA}}$ token A in one of the smallest pools with index $i^{\ast}$, where $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the trader has no less than the revenue of any other actions:

### v2

@@ VIII-A Simulation Setup
(a) Relative traders’ costs to the external market.
(b) Relative liquidity providers’ revenue to Uniswap.
VIII-B Trade Distribution and Splits
@@
We want to confirm that the distribution of trades in SAMM is balanced and that trade splits are infrequent. Figure 6 illustrates the number of trades executed in each shard compared to the average. Error bars show the range of relative trade numbers in each shard compared to the average. The difference from the average is always under $5\%$. Additionally, we analyze the proportion of trades that are split into multiple transactions within SAMM. Figure 7 shows that with $c=0.01$ trade splits are infrequent, occurring in less than $8\%$ of trades, even with 32 shards. Furthermore, the total number of trade splits results in less than a $45\%$ increase in overall trade volume, a relatively minor increment given the five-fold throughput improvement. When the number of shards is larger or $c$ is smaller, the proportion of trade splits is higher.
This confirms that our theoretical analysis matches actual trading behavior. In particular, SAMM incentivizes parallel execution can be ensured without relying on theoretical assumptions such as the presence of perfect arbitrageurs.
VIII-C Cost and Revenue
We compare the revenue and cost of liquidity providers and traders, respectively. Figure 8(a) shows the average ratio of traders’ costs in SAMM compared to the costs in the external market, considering different numbers of shards and varying values of $c$. We observe that with larger $c$, the cost for traders increases, as expected. For a fixed $c$, the cost for traders decreases as the number of shards increases, aligning with SAMM’s $c$-smaller-better property, where more shards result in less liquidity in each shard. Additionally, we compare the cost for traders in Uniswap, depicted by the black line. In all cases, the cost of SAMM is either smaller ($c=0.003$ with no less than 2 shards) or slightly larger than that of Uniswap, differing by less than $1\%$ with $c=0.01$ and $0.3\%$ with $c=0.005$, when there are at least 2 shards. We aggregate trading fees across all shards to evaluate the revenue of liquidity providers. Figure 8(b) shows SAMM consistently generates higher revenue than Uniswap, with $c=0.01$ and 7 shards yielding over 15 times the revenue. Initially, revenue in SAMM rises with more shards due to increased throughput but with even more shards it declines as trading fees per trade decrease in smaller shards (see Appendix G, Lemma G-B1).
@@ VIII-C Cost and Revenue
SAMM significantly outperforms Uniswap in terms of liquidity provider revenue while maintaining comparable costs for traders. This increase without major cost hikes is due to SAMM’s enhanced throughput, allowing for more trades and higher total trading fees.
VIII-D Throughput and Incentive Trade-off
Note that the results above imply a trade-off between participant incentives and system performance. When $c$ is larger, there are fewer trade splits and higher revenue for liquidity providers, but the cost for traders increases. Conversely, fixing $c$ and increasing the number of shards results in lower trader costs but more trade splits, which can negatively impact system performance. Additionally, liquidity provider revenue peaks at a certain number of shards. Therefore, it is crucial for system designers to choose an appropriate $c$ and number of shards to balance participant incentives and system performance.
IX Conclusion
@@
We present SAMM, a scalable AMM. The key enabler of SAMM is the design of a trading fee mechanism that incentivizes parallel operations. We analyze trader and liquidity provider behaviors as a game, showing that parallel operations are the best response, and validate by simulation with real trade traces. We implement and deploy SAMM in local testnets of Sui and Solana, demonstrating more than 5x and 16x throughput improvement, up to the underlying system’s limits. Our results indicate that reducing serial bottlenecks of independent contracts should be a focus of smart-contract platforms to allow for AMM scaling (See Appendix H). Meanwhile, SAMM can be directly deployed to scale AMMs on existing platforms, for direct use and as part of the DeFi eco-system.
@@ Appendix A Growing Demand for AMM
We analyze the demand for AMMs by calculating the number of trades per second on the prominent Uniswap versions 1-3, from its deployment in November 2018 to March 2024. We use data from Dune [^3] https://dune.com/, a comprehensive database for blockchain data. Figure 9 illustrates the exponentially increasing demand, from $0.78$ average trades per second in 2020 to $9.54$ in 2024. On the Ethereum blockchain, most Uniswap transactions are executed through versions 2 and 3, incurring gas costs of $152.809$ and $184.523$ respectively. Given that the average total gas per block is $15,000,000$ and the block interval is $12$ seconds, Ethereum can facilitate up to $8.18$ trades per second for v2 and $6.77$ trades per second for v3. However, in March 2024, the average monthly trades on Uniswap reached $13.9$ per second, nearly doubling Ethereum’s processing capacity. Therefore, most demand of Uniswap is processed through off-chain solutions [2].
@@ Appendix C CPMM Does Not Satisfy Either property
{thm}
[] For any value of $10<c<$ 1, the CPMM cost function $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$ does not satisfy either the $c$-Non-Splitting or the $c$-smaller-better properties.
Proof:
@@
Appendix D Proof of Proposition V-C
  • Proposition V-C (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@ Appendix D Proof of Proposition V-C
The $c$-smaller-better property considers two pools $i$ and $j$, assuming their reported prices are identical. Denote the inverse of this price by $c^{\textit{AB}}$, so $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}=\frac{1}{c^{\textit{AB}}}>0$. The property requirement (Equation V-A2) becomes
@@
If the polynomial value is greater than the bound, $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}>r_{\max}$, then from Equations 49, and 10, the gross amount becomes
@@
Combining the constraints we just found (Equations 31 and 32) into the gross amount expression (Equations 9 and 10) we have
@@
From Equations 27 and 32, we have
@@
Appendix E Proof of Theorem V-C
  • Theorem V-C (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@ Appendix E Proof of Theorem V-C
Then we can expand the trading fee function and the gross amount function to:
@@
$\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{B}}{(1-c)^{3}\left(R^{A}\right)^{2}}\leq 0.$
The above inequality is equal to
@@
The above condition is sufficient for the $c$-Non-Splitting property.
@@
Therefore, Item 1 is sufficient for Equation 50, which indicates that the $c$-Non-Splitting property holds under Item 1.
@@
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}=\\ -c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}\,\,.$
@@
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}>\\ -c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-cR^{A})^{2}}\,\,.$ (53)
@@
The above equation is equivalent to
@@ Appendix F CPMM Equilibrium
In CPMM, the gross amount is linear to the net amount. Since traders can suffer less from slippage by splitting a trade, the gross amount of splitting a trade is less than the gross amount of trading the same amount at one time. {thm}[] In $\Gamma_{n}(\textit{tf}_{\textit{CPMM}})$, the following strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}})$ is the only dominant strategy for the BA trader, where
@@
Proof:
We start by calculating the best response for the BA trader. Since the utility $U^{BA}(\textbf{R},b^{\textit{BA}},\pi^{BA})$ is a linear combination of the revenue over actions $U^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ (Equation 15), we can first calculate the optimal action for the BA trader. Combining the CPMM gross amount (Equation IV-C) and the revenue function (Equation VI-A4), we obtain
@@
We first consider the case of two pools and then extend it to the general case. We define the function $f(\cdot,\cdot)$ which is proportional to the utility of the trader in $\textit{pool}_{i}$ and $\textit{pool}_{j}$:
$f(b^{\textit{BA}}_{i},b^{\textit{BA}}_{j})\coloneqq-\left(\frac{R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}+\frac{R^{A}_{j}b^{\textit{BA}}_{j}}{R^{A}_{j}-b^{\textit{BA}}_{j}}\right)\,\,,$
. Denote by $z\coloneqq b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j}\leq b^{\textit{BA}}$. Then, we have
@@
$\tau^{BA}(\textbf{R},b^{\textit{BA}})=\begin{cases}1,&\text{if }a^{\textit{BA}}=\left(\frac{R^{A}_{1}}{\sum R^{A}_{i}},\cdots,\frac{R^{A}_{n}}{\sum R^{A}_{i}}\right)\\ 0,&\text{Otherwise.}\end{cases}$ (60)
Appendix G SAMM Equilibrium
We analyze the behavior of players in the game with SAMM. We first prove that the trader randomly selects a pool to trade when the states of pools are balanced (§G-A). Then, we show the system tends to the balanced state since liquidity providers invest their tokens in the smallest pools, reducing the difference in the volume of pools (§G-B).
G-A Trader Strategy
Consider the case that the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. As discussed in Section V-C, the SAMM gross amount satisfies the $c$-non-splitting property and $c$-smaller-better property for a certain $10<c<$ 1. We assume that the required amount of token A, $b^{\textit{BA}}$, is at most a fraction $c$ of the amount of deposited token A in all pools, i.e.,
$\forall 1\leq i\leq n,b^{\textit{BA}}\leq cR^{A}_{i}\,\,.$
G-A1 Traders’ optimal action
The $c$-non-splitting property and $c$-smaller-better property give a trader the incentive to randomly select one of the smallest pools to trade all her required tokens. Recall that $a_{i}^{\text{BA}}(b^{\textit{BA}})$ is the action of acquiring all $b^{\textit{BA}}$ token A in $\textit{pool}_{i}$ (Equation 20). We define the set of actions that trade in one of the smallest pools: {defi}[] The Smallest Pool Action Set is the set of actions that acquire all $b^{\textit{BA}}$ token A in one of the smallest pools under state R:
{lem}
[] In $\Gamma_{n}(tf_{\textit{SAMM}})$, a trader wants to get $b^{\textit{BA}}$ token A when the system state is R. Then for the action which obtains all $b^{\textit{BA}}$ token A in one of the smallest pools with index $i^{\ast}$, where $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the trader has no less than the revenue of any other actions: